# StreetCrafter: Street View Synthesis with Controllable Video Diffusion Models

### v2

@@ 3.3 Crafting Dynamic 3D Scenes
To better align the generated video frames from the diffusion model with the 3DGS scene representation, we generate samples from the noisy rendered latent instead of pure noise inspired by previous works [63, 43]. We find this can help maintain the overall scene structure and accelerate the training process due to fewer denoising steps. To be specific, we render images $\{{\mathbf{I}}_{i}^{r}\}_{i=1}^{K}$ from novel views $\{\hat{\mathbf{C}}_{i}\}_{i=1}^{K}$, which are encoded and perturbed to noisy latents $\{\mathbf{z}_{i,t}^{r}\}_{i=1}^{K}$ given diffusion timestep $t$ derived from noise scale $s$. Then we generate samples from the denoising U-Net by running the DDIM sampling [46]. The samples are further decoded to novel view images $\{\hat{\mathbf{I}}_{i}\}_{i=1}^{K}$ for supervision. Based on the following generation process, we employ a progressive optimization strategy by gradually reducing the noise scale $s$. This helps the model learn to remove artifacts by relying more on the diffusion prior in the early training stage and progressively focusing on refining details as the training advances.
@@ 4 Implementation Details
We initialize StreetCrafter from the pretrained checkpoint of Vista [12]. We first train all the parameters of the video denoising U-Net at the resolution of 320 $\times$ 576 with batch size of 16 and learning rate of $5e^{-5}$ for 30000 iterations. Then we fix the temporal layers and finetune the spatial layers of U-Net at the resolution of 576 $\times$ 1024 with batch size of 8 and learning rate of $1e^{-5}$ for another 3000 iterations. During training, we randomly drop the reference image and LiDAR conditions with a probability of 15% independently. It takes 2 days to train the model on 8 NVIDIA A800 80GB GPUs using Adam [26] optimizer. During inference, we set the sampling steps to 50 and classifier-free guidance (CFG) [17] scale to 2.5 and generate videos of length $n=25$ at the resolution of $576\times 1024$. For novel trajectories longer than $n$, we iteratively sample $n$-length video frames with an overlapping frame length of 5 to construct the full-length video. The temporal window size $l$ is set to cover the LiDAR point cloud from adjacent ±$\pm 1s$ frames.
@@ 8.1 More Comparisons
Comparisons with baselines.  We provide more qualitative comparisons on Waymo [48] dataset under the setting of lane change in Figure 13. Figures 14, 15 display the view interpolation results on Waymo Open [48] and PandaSet [65] datasets. Our method achieves comparable rendering quality to the baselines while achieving significantly better results for view extrapolation.
@@ 8.3 More Ablations
Figure 11: More editing results on the Waymo [48] dataset. Images in the right and left columns represent the results before and after editing, respectively.
@@
We conduct experiment on one Waymo sequence to analyze the influence of noise scale $s$.s We have demonstrated in the main paper that adding noise to the render latents leads to better scene consistency than starting from gaussian noise. Since the added noise would have little influence when $s$ is less than 0.3 according to the sampling scheme of Vista [24], we set $s_{\text{min}}$ to 0.3 and ablate on the value of $s_{\text{max}}$. The results in Table 7 indicates that reducing $s$ from 0.7 to 0.3 maintains a balance between sampling steps and rendering quality.
@@ 8.4 Deformable Objects
We show the generated videos of StreetCrafter under scene with multiple deformable objects such as pedestrians in Figure 16. Although multi-frame LiDAR conditions are not ideal for deformable objects, our model can generate plausible results with the generative prior of diffusion model.
@@
Figure 13: Qualitative comparisons on the Waymo [48] dataset. The camera is laterally shifted for 2 meters to left or right. Input view refers to the closest training camera.

### v3

@@ 3.3 Crafting Dynamic 3D Scenes
To better align the generated video frames from the diffusion model with the 3DGS scene representation, we generate samples from the noisy rendered latent instead of pure noise inspired by previous works [63, 43]. We find this can help maintain the overall scene structure and accelerate the training process due to fewer denoising steps. To be specific, we render images $\{{\mathbf{I}}_{i}^{r}\}_{i=1}^{K}$ from novel views $\{\hat{\mathbf{C}}_{i}\}_{i=1}^{K}$, which are encoded and perturbed to noisy latents $\{\mathbf{z}_{i,t}^{r}\}_{i=1}^{K}$ given diffusion timestep $t$ derived from noise scale $s$. Then we iteratively denoise the latents to clean samples by running the EDM sampling [24] used in Vista. The samples are further decoded to novel view images $\{\hat{\mathbf{I}}_{i}\}_{i=1}^{K}$ for supervision. Based on the following generation process, we employ a progressive optimization strategy by gradually reducing the noise scale $s$. This helps the model learn to remove artifacts by relying more on the diffusion prior in the early training stage and progressively focusing on refining details as the training advances.
@@ 4 Implementation Details
We initialize StreetCrafter from the pretrained checkpoint of Vista [12]. We first train all the parameters of the video denoising U-Net at the resolution of 320 $\times$ 576 with batch size of 16 and learning rate of $5e^{-5}$ for 30000 iterations. Then we fix the temporal layers and finetune the spatial layers of U-Net at the resolution of 576 $\times$ 1024 with batch size of 8 and learning rate of $1e^{-5}$ for another 3000 iterations. During training, we randomly drop the reference image and LiDAR conditions with a probability of 15% independently. It takes 2 days to train the model on 8 NVIDIA A800 80GB GPUs using Adam [26] optimizer. During inference, we set the sampling steps to 50 and classifier-free guidance (CFG) [17] scale to 2.5 and generate videos of length $n=25$ at the resolution of $576\times 1024$. For novel trajectories longer than $n$, we iteratively sample $n$-length video frames with an overlapping frame length of 5 to construct the full-length video. The temporal window size $l$ is set to cover the LiDAR point cloud from adjacent $\pm 1s$ frames.
@@ 8.1 More Comparisons
Comparisons with baselines.  We provide more qualitative comparisons on Waymo [47] dataset under the setting of lane change in Figure 13. Figures 15, 16 display the view interpolation results on Waymo Open [47] and PandaSet [65] datasets. Our method achieves comparable rendering quality to the baselines under input trajectory while achieving significantly better results for view extrapolation.
@@
8.2 Adapting to Advanced Video Diffusion Model
We can further enhance the performance of StreetCrafter by adapting it to more advanced video diffusion models, such as Wan [56]. To be specific, we adopt Wan2.1-14B-I2V as our base model and patchify the video and LiDAR condition latents separately with the same embedding layer. The LiDAR condition latents are first passed through a zero-initialized linear layer and then element-wise added to the video latents before being fed into the diffusion transformer blocks. During training, instead of optimizing all the parameters as in Vista, we utilize a Low-Rank Adaptation approach (LoRA) [22] for the diffusion transformer while keeping the pretrained parameters frozen to improve the training efficiency and preserve the 2D video prior.
As shown in Figure 11, StreetCrafter based on Wan can generate high quality results even under extreme viewpoint changes. For instance, the regions highlighted in Figure 11 are completely invisible to the input camera and LiDAR sensors, which means no condition is available for these areas during inference. Nevertheless, our method can still produce plausible results thanks to the powerful generative prior of the base model. This demonstrates a key strength of our method: it can be readily adapted to more advanced video diffusion models, achieving improved performance with minimal effort.
@@ 8.3 More Editings
Figure 11: Visual results of StreetCrafter based on Wan2.1 [56]. When having a more powerful video generation model, StreetCrafter is able to produce realistic results under challenging scenes with severe occlusions and diverse weather conditions. The camera is gradually shifted for 3 meters to left or right.
@@ 8.4 More Ablations
We conduct experiment on one Waymo sequence to analyze the influence of noise scale $s$. We have demonstrated in the main paper that adding noise to the render latents leads to better scene consistency than starting from gaussian noise. Since the added noise would have little influence when $s$ is less than 0.3 according to the sampling scheme of Vista [24], we set $s_{\text{min}}$ to 0.3 and ablate on the value of $s_{\text{max}}$. The results in Table 7 indicates that reducing $s$ from 0.7 to 0.3 maintains a balance between sampling steps and rendering quality.
@@ 8.5 Deformable Objects
We show the generated videos of StreetCrafter under scenes with multiple deformable objects such as pedestrians in Figure 17. Although multi-frame LiDAR aggregation leads to incorrect guidance for deformable objects, our method can produce plausible results thanks to the generative prior of video diffusion model.
@@
Figure 13: Qualitative comparisons on the Waymo [47] dataset. The camera is gradually laterally shifted for 3 meters to left or right.
Figure 14: More editing results on the Waymo [47] dataset. The right and left images represent the results before and after editing.